# Patrick Kahardipraja1, Thomas Wiegand1,2,3, Wojciech Samek1,2,3, Sebastian Lapuschkin1

### v1

@@ E.1 Quantitative Analysis
This becomes particularly evident when comparing MaxAct* to feature descriptions generated for Llama-3.1-8B-Instruct in Choi et al. (2024). While clarity scores are comparable—albeit slightly lower for MaxAct* — responsiveness and purity show significant improvements. This difference may partially stem from dataset construction variations. Notably, the purity distributions of the two approaches are strikingly different, even opposing: MaxAct* exhibits a right-skewed peak, whereas Transluce’s feature descriptions perform poorly overall. Faithfulness differences are minor but still favor MaxAct*, likely due to the generation of higher-quality feature descriptions.
@@ E.2 Qualitative Analysis
Fine-tuning automated interpretability requires consideration of all metrics in [Uncaptioned image]FADE. The comparison of metric distributions against the simulated activation-based metric in Figure 3 highlights that relying solely on this metric is insufficient for accurately assessing the quality of feature descriptions. If an automated interpretability framework is fine-tuned exclusively on such a metric, it may generate suboptimal descriptions.
@@
SAEs descriptions often fail to accurately capture the concept. Figure 10 presents the heatmap for feature 475, whose description emphasizes the occurrence of the token “self.” However, the feature does not activate on all instances of “self” (highlighted in yellow), indicating that a crucial aspect of the concept is missing or remains unclear. Additionally, the feature activates on other tokens, further suggesting that the description is incomplete.
This is reflected in the evaluation metrics: low clarity indicates that the concept is not expressed precisely enough for the evaluating LLM to generate synthetic data that reliably activates the feature. A responsiveness score of 0.93 suggests that the feature does activate on natural data aligned with the concept, while a purity score of 0.81 reveals that although the feature is primarily associated with the described concept, it also responds to other inputs.
Similar issues arise in features 653 and 821, where the underlying concepts appear highly specific – activating on a particular token within a specific context. However, their descriptions are overly broad, making it difficult to generate synthetic data that reliably triggers the feature. These and many other similar features contribute to the left-skewed peak in the clarity distribution, which becomes more pronounced for narrower concepts represented by features in Gemma Scope 65K.
@@
Out-of-Distribution Features in Gemma Scope 65K SAEs. The dataset used for automatic interpretability omitted certain concepts, such as "new line," leading to gaps in feature descriptions. These omissions contribute to the small left-side peak in responsiveness distribution in Figure 5. Several features, including 3315, 3858, and 4337, lack activation heatmaps under the MaxAct* approach, as the dataset does not represent their concepts. Consequently, the explainer model, relying on unrelated sentences, generates incorrect descriptions (see Figure 12). Heatmaps from Neuronpedia [^2] https://www.neuronpedia.org/gemma-2-2b/20-gemmascope-res-65k/3286 reveal what would activate these features, highlighting limitations of the dataset, used in this work, and broader issues in the automated interpretability pipeline. For example, despite obtaining and visualizing correct results, feature descriptions available on Neuronpedia are also not representing a correct concept. Similar results have been obtained for the <bos> token and indentation in text and code.
@@
Reliable Evaluation — [Uncaptioned image]FADE Identifies the Best Description. Different automated interpretability methods prioritize either activation-based metrics or faithfulness-based measures, leading to descriptions that may be overly broad or inaccurate.

### v2

@@ C R P F C R P F C R P F C R P F
This becomes particularly evident when comparing MaxAct* to feature descriptions generated for Llama-3.1-8B-Instruct in Choi et al. (2024). While Clarity scores are comparable – albeit slightly lower for MaxAct* – Responsiveness and Purity show significant improvements. This difference may partially stem from the fact, that the dataset used in this work is significantly larger than the one used to produce feature descriptions in Choi et al. (2024). Notably, the Purity distributions of the two approaches are strikingly different, even opposing: MaxAct* exhibits a right-skewed peak, whereas Transluce’s feature descriptions perform poorly on this metric overall. Faithfulness differences are minor but still favor MaxAct*, likely due to the generation of higher-quality feature descriptions.
Metrics correlation depends on the quality of produced feature descriptions
In general, we observe a strong correlation between the evaluated metrics, with the exception of the Faithfulness metric. Importantly, the degree of correlation is influenced by the quality of the feature descriptions being assessed (see Table 16). Descriptions generated using MaxAct* exhibit the strongest alignment with their corresponding features, whereas Neuronpedia descriptions, especially on Gemma Scope 65K, show the weakest alignment, which results in higher correlation between Clarity, Purity and Responsiveness metrics for MaxAct* Gemma Scope, and lower correlation between Faithfulness and other metrics.
This trend can be explained by the fact that well-aligned descriptions tend to yield consistently high values across all metrics. Conversely, when the descriptions are of lower quality, the correlation decreases, as each metric is sensitive to different types of failure cases (see Appendix E.2).
An exception to this pattern is the Faithfulness metric. In its case, the relationship is reversed: poorer description quality leads to a higher correlation with the other metrics. This occurs because Faithfulness is set to zero when other metrics are low and the Faithfulness score is not being computed.
@@ E.2 Qualitative Analysis
Fine-tuning automated interpretability requires consideration of all metrics in [Uncaptioned image]FADE
The comparison of metric distributions against the simulated activation-based metric in Figure 3 highlights that relying solely on this metric is insufficient for accurately assessing the quality of feature descriptions. If an automated interpretability framework is fine-tuned exclusively on such a metric, it may generate suboptimal descriptions.
@@
Out-of-Distribution Features in Gemma Scope 65K SAEs
The dataset used for automatic interpretability omitted certain concepts, such as "new line," leading to gaps in feature descriptions. These omissions contribute to the small left-side peak in Responsiveness distribution in Figure 6. Several features, including 3315, 3858, and 4337, lack activation heatmaps under the MaxAct* approach, as the dataset does not represent their concepts. Consequently, the Explainer Model, relying on unrelated sentences, generates incorrect descriptions (see Figure 13). Heatmaps from Neuronpedia [^2] https://www.neuronpedia.org/gemma-2-2b/20-gemmascope-res-65k/3286 reveal what would activate these features, highlighting limitations of the dataset used in this work, and broader issues in the automated interpretability pipeline. For example, despite obtaining and visualizing correct results, feature descriptions available on Neuronpedia are also not representing a correct concept. Similar results have been obtained for the <bos> token and indentation in text and code.
@@
Reliable Evaluation – [Uncaptioned image]FADE Identifies the Best Description
Different automated interpretability methods prioritize either activation-based metrics or Faithfulness-based measures, leading to descriptions that may be overly broad or inaccurate.